OpenStudy (anonymous):

I'll give a medal One positive number is 5 times another number. The difference between the two numbers is 556, find the numbers. The two numbers in increasing order are

OpenStudy (yoloshroom):

Hmm let me work this one out

OpenStudy (anonymous):

5x=y y-x=556 +x +x y=x+556 5x=x+556 /5 /5 x=1/5x+111and1/5 -1/5x 4/5x=111and1/5 /4/5 /4/5 x=139 139+556=695=y

OpenStudy (anonymous):

X x 5y=556

OpenStudy (anonymous):

IDK Im tired Ima go to science Bye :)